4 times Randy's age decreased by 3 is equal to 3 times Randy's age increased by7 How old is Randy
Answer by tifftiff133(4)

To solve this problem you should start by writing it out. Randy's age- R 4*R-3=3*R+7 Now add or subtract following order of operations. Take 3 over through addition. 4*R=3*R+7 Then take the 3*R (3R) over to the other side using subtraction. 1R=10 Randy's age is 10. You should probably check my work because I do make mistakes, but that is what I got. I hope I was at least a little helpful. Good Luck on your math work!